an 80.0 kg athlete running a "4-minute mile" (i.e. 4.00 min/mile)

nm =?
1
Expert's answer
2019-12-11T07:54:35-0500

1 mile / 4 min x (1 min / 60 s) x (1609 m / 1 mile) = 6.70 m/s

wavelength = 6.626x10-34J*s/(80kg*6.7m/s) = 3.55*10-31m = 3.55*10-22 nm